OpenStudy (anonymous):

how do I solve this for x? 17x+x^2= 90

OpenStudy (accessdenied):

Do you know how to solve the equation in this form: \(ax^2 + bx + c = 0\) ?

OpenStudy (anonymous):

kinda?? I really can't remember....

OpenStudy (accessdenied):

We would want to put it into that form. There are ways we can find the solutions when all the information is on one side. :) So, we just have to subtract off the 90 from both sides and we should have this: \( x^2 + 17x - 90 = 0\)

OpenStudy (anonymous):

ok, then where do we go?

OpenStudy (accessdenied):

It doesn't seem factorable to me, so we have to use something else. Do you recall the quadratic formula? \[ \quad x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \] Basically, we just have to plug in \(a = 1\), \(b=17\), and \(c = -90\) here and it should give us the solutions. :)

OpenStudy (anonymous):

oh! the quadratic formula!:) i remember now! thanks so much for your help!!

OpenStudy (accessdenied):

You're welcome. :)

OpenStudy (accessdenied):

I get: -17 +- sqrt(649) ----------------- 2 649 seems to only break into 11 and 59 as prime factorization, so nothing really comes out of the radical.

OpenStudy (anonymous):

yea i got it:)